In a triangleabc anglea is an obtuse angle and sine a is 3/5 and cost is 12/13 then find since?

Grade:11

1 Answers

Arun
6 years ago
sinA = 3/5 so cosA = - 4/5 (since A obtuse); similarly, using a simple right-angled triangle if cosB = 12/13, then sinB = 5/13. C = 180 - (A + B) so sinC = sin[180 - (A + B)] = sin(A + B) (4 quadrants) so sinC = sinAcosB + cosAsinB = (3/5)*(12/13) + (-4/5)*(5/13) = 36/65 - 20/65 = 16/65